OpenStudy (anonymous):

Find a if b = 60 and c = 61. A. a = 121 B. a = 1 C. a = 11 D. a = √7321

OpenStudy (mathstudent55):

Is this a right triangle? a^2 + b^2 = c^2 a^2 + 60^2 = 61^2 a^2 = 61^2 - 60^2 a = sqrt(61^2 - 60^2)

OpenStudy (anonymous):

i dont know if that changes things?

OpenStudy (mathstudent55):

It confirms that I understood it even without the picture, so just follow what I did above.

OpenStudy (hoopcoach99):

SO is the answer 121

OpenStudy (mathstudent55):

a^2 = 121 a = 11 The answer is a = 11.